# DSL or Code? Evaluating the Quality of LLM-Generated Algebraic Specifications: A Case Study in Optimization at Kinaxis

### v1

@@ 5.8. Validity Considerations
Reliability. We release the EXEOS code, prompts, and the Public dataset (Ayoughi, 2025), but cannot share the Industry dataset. This reduces pretraining-leakage risk but limits replication. To improve reliability, we document our experimental procedure in detail.
@@ 7. Related Work
Automated assistance for extracting models from text has been studied prior to LLMs, e.g., (Yue et al., 2011; Arora et al., 2016, 2019). With LLMs, this research has accelerated, producing more effective methods for generating various models, including goal models (Chen et al., 2023a), domain models (Yang et al., 2024; Silva et al., 2024; Chen et al., 2025), sequence diagrams (Ferrari et al., 2024; Jahan et al., 2024), and activity diagrams (Khamsepour et al., 2025; Chen et al., 2025). Wang et al. (Wang et al., 2023) propose grammar prompting to improve few-shot DSL generation by pairing each example with a small grammar fragment that captures the relevant syntax rules. The LLM first predicts such a grammar for the input and then generates the output under this grammar’s rules. In contrast, EXEOS does not impose a formal grammar during generation. Instead, it relies on few-shot examples and domain-specific syntax guidelines, and it introduces a structuring step that identifies problem components prior to generating AMPL models. Recently, Joel et al. (Joel et al., 2025) have conducted a systematic review of LLM-based code generation for DSLs. Their findings identify iterative feedback from external tools, such as compilers or solvers, as a technique that improves LLM accuracy; this aligns with the use of solver diagnostics in our work.
LLMs have also been used to generate logic-based artifacts, such as temporal-logic formulas (Cosler et al., 2023; Chen et al., 2023b), OCL constraints (Abukhalaf et al., 2024), and assertion-based postconditions (Endres et al., 2024). These artifacts can be validated only in combination with a host artifact (e.g., a formal model for temporal logics, Java code for postconditions, or a UML model for OCL), making the effectiveness of their automated derivation inseparable from that context and thus precluding direct compiler- or solver-style diagnostics. Closer to our setting, work on generating optimization specifications (AhmadiTeshnizi et al., 2024; Zhang et al., 2024) does yield solver-executable artifacts with solver feedback driving refinement. However, these approaches do not explicate how optimization data is acquired and bound, and couple such tasks with specification extraction, making them ill-suited for industrial-scale datasets. In contrast, informed by practice, EXEOS introduces a dedicated data-handling step, decoupled from specification extraction, to ensure reliable data binding at scale. Finally, to our knowledge, no prior work on extracting optimization specifications from text provides a systematic comparison of LLMs in DSLs versus general-purpose languages, as we do, or ablates components to quantify their impact.

### v2

@@ 5.8. Validity Considerations
Reliability. We release the EXEOS code, prompts, and the Public dataset (neayoughi2025optigen), but cannot share the Industry dataset. This reduces pretraining-leakage risk but limits replication. To improve reliability, we document our experimental procedure in detail.
@@ 7. Related Work
Automated assistance for extracting models from text has been studied prior to LLMs, e.g., (Yue11; AroraSBZ16; AroraSNB19). With LLMs, this research has accelerated, producing more effective methods for generating various models, including goal models (Chen23), domain models (YangCCMV24; silva2024application; Chen25), sequence diagrams (FerrariAA24; JahanHGRRRS24), and activity diagrams (khamsepour2025impact; Chen25). Wang et al. (WangW0CSK23) propose grammar prompting to improve few-shot DSL generation by pairing each example with a small grammar fragment that captures the relevant syntax rules. The LLM first predicts such a grammar for the input and then generates the output under this grammar’s rules. In contrast, EXEOS does not impose a formal grammar during generation. Instead, it relies on few-shot examples and domain-specific syntax guidelines, and it introduces a structuring step that identifies problem components prior to generating AMPL models. Recently, Joel et al. (JoelWF25) have conducted a systematic review of LLM-based code generation for DSLs. Their findings identify iterative feedback from external tools, such as compilers or solvers, as a technique that improves LLM accuracy; this aligns with the use of solver diagnostics in our work.
LLMs have also been used to generate logic-based artifacts, such as temporal-logic formulas (CoslerHMST23; ChenGZF23), OCL constraints (AbukhalafHK24), and assertion-based postconditions (EndresFCL24). These artifacts can be validated only in combination with a host artifact (e.g., a formal model for temporal logics, Java code for postconditions, or a UML model for OCL), making the effectiveness of their automated derivation inseparable from that context and thus precluding direct compiler- or solver-style diagnostics. Closer to our setting, work on generating optimization specifications (AhmadiTeshniziG24; ZhangWGWLYY24) does yield solver-executable artifacts with solver feedback driving refinement. However, these approaches do not explicate how optimization data is acquired and bound, and couple such tasks with specification extraction, making them ill-suited for industrial-scale datasets. In contrast, informed by practice, EXEOS introduces a dedicated data-handling step, decoupled from specification extraction, to ensure reliable data binding at scale. Finally, to our knowledge, no prior work on extracting optimization specifications from text provides a systematic comparison of LLMs in DSLs versus general-purpose languages, as we do, or ablates components to quantify their impact.